# A Flexible, Equivariant Framework for Subgraph GNNs via Graph Products and Graph Coarsening

### v1

@@ 4.1 Product graph definition
$A^{\mathcal{T}}(S_{1},S_{2})=\begin{cases}1&\text{if }\exists v\in S_{1},\exists v\in S_{2}\text{ s.t. }A(v,u)=1,\\ 0&\text{otherwise},\end{cases}$ (3)
To clarify, in our running example (Figure 1), it holds that $A^{\mathcal{T}}(\{a,b,c,d\},\{e\})=1$, while $A^{\mathcal{T}}(\{e\},\{f\})=0$. For a more formal definition, refer to Definition A.3.
Our implementation of the graph coarsening function $\mathcal{T}$ employs spectral clustering [34] [^5] Other graph coarsening or clustering algorithms can be readily used as well. to produce $T$ clusters, which in our framework controls the size of the bag.
Defining $\mathcal{T}(G)\square G$.
@@
The connectivity in this product graph induces the horizontal ($\mathcal{A}_{G}$) and vertical updates ($\mathcal{A}_{\mathcal{T}(G)}$) in Equation 2, visualized in Figure 1(right) via grey and red redges, respectively.
@@ 4.2 Symmetry-based updates
In the previous subsection, we relied on a combination of a coarsening function and the graph Cartesian product to derive the two induced connectivities $\mathcal{A}_{G},\mathcal{A}_{\mathcal{T}(G)}$ of our product graph. As mentioned before, these connectivities are used to perform message-passing on our product graph (see Equation 2).
Inspired by recent literature on Subgraph GNNs [12, 3, 39], which incorporated and analyzed additional non-local updates arising from various symmetries (e.g., updating a node’s representation via all nodes in its subgraphs), this section aims to identify potential new updates that can be utilized over our product graph. More specifically, in this section we focus on the third term in Equation 2, dubbed Symmetry-based updates. To that end, we study the symmetry structure of the node feature tensor in our product graph, $\mathcal{X}(S,v)$. For better clarity in this derivation, we change the notation from nodes ($v$) to indices ($i$).
@@ 4.2.1 Symmetries of our product graph
where we define the action of $\sigma\in S_{n}$ on a set $S=\{i_{1},i_{2},\ldots,i_{k}\}$ of size $k$ as: $\sigma\cdot S\triangleq\{\sigma^{-1}(i_{1}),\sigma^{-1}(i_{2}),\ldots,\sigma^{-1}(i_{k})\}\triangleq\sigma^{-1}(S)$.
@@ 4.2.2 Derivation of linear equivariant layers for the node feature tensor
Here we will characterize the linear equivariant layers with respect to the symmetry defined above, focusing on Equation 6. We adopt a similar notation to [22], and note that, in our derivation, we assume for simplicity that the number of feature channels is $d=1$ (extension to multiple features is straightforward [22]). In addition, our analysis considers the more general case where $V^{\mathcal{T}}$ encompasses all potential super-nodes formed by subsets of $[n]$.
Our main tool is the characterization of linear equivariant layers for permutation symmetries as parameter-sharing schemes [35, 31, 22]. In a nutshell, this characterization states that the parameter vectors of the biases, invariant layers, and equivariant layers can be expressed as a learned weighted sum of basis tensors, where the basis tensors are indicators of the orbits induced by the group action on the respective index spaces. We focus here on presenting the final results; detailed discussion and derivations are available in Appendix E. Our main results are summarized in Proposition E.1.
Equivariant bias and invariant layers. The bias vectors as well as the weight vectors defining linear layers in our space are in $\mathbb{R}^{2^{n}\times n}$. Hence, we define an equivalence relation $\sim$ in the associated index space $\{1,\dots,2^{n}\times n\}$ which, as stated before, corresponds to the orbits under the action of $S_{n}$. For pairs $(S,i)\in(\mathcal{P}([n])\times[n])$, $\gamma^{k^{+}}$ corresponds to all pairs with $|S|=k$ and $i\notin S$, and $\gamma^{k^{-}}$ to all pairs with $|S|=k$ and $i\in S$, leading to the following partition:
This partition effectively splits the indices $(S,i)$ according to the size of $S$ and whether $i\in S$. Given $\gamma\in(\mathcal{P}([n])\times[n])/_{\sim}$, the following is a basis tensor of the bias vector of the invariant layer,
@@
Weight matrices. Following similar reasoning, consider elements $(S_{1},i_{1},S_{2},i_{2})\in(\mathcal{P}([n])\times[n]\times\mathcal{P}([n])\times[n])$. To characterize the orbits of $S_{n}$ in this space we define partitions according to six conditions. For example, two of these conditions are the sizes of $S_{1}$ and $S_{2}$, which remain invariant under permutations. Another condition is the size of the intersection set $S_{1}\cap S_{2}$, which also remains invariant under permutations. For a thorough discussion and derivation of the six conditions, please refer to Appendix E. We define $\sim$ as the equivalence relation for these six conditions. Given an equivalence class, or orbit, $\Gamma\in(\mathcal{P}([n])\times[n]\times\mathcal{P}([n])\times[n])/_{\sim}$, we define a basis tensor, $\mathbf{B}^{\Gamma}\in\mathbb{R}^{2^{n}\times n\times 2^{n}\times n}$ by setting:
@@
A visualization of the two basis vectors, Equations 8 and 9, is available in Figure 3. The following (informal) proposition summarizes the results in this section (the proof is given in Appendix G),
@@ 4.2.3 Incorporating symmetry-based updates in our framework
To begin, we note that any linear equivariant layer
@@
can be realized through an MPNN  [13] applied to a fully connected graph with appropriate edge features. This is formally stated in Lemma F.1, the main idea is to encode the parameters on the edges of this graph (see visualization inset). Thus, the natural construction of $\mathcal{A}_{\text{Equiv}}$ corresponds to a fully connected graph, with appropriate edge features derived from the parameter-sharing scheme we have developed.
@@
However, one of our main goals and guidelines in developing our flexible framework is to maintain efficiency, and to align with the maximally expressive GNN, namely GNN-SSWL$+$ [39, 3], for the case of a trivial coarsening function, $\mathcal{T}(G)=G$ (which correspond to the full-bag setting). To achieve this, we opt for a sparser choice by using only a subset of the basis vectors (defined in Equation 9) to construct $\mathcal{A}_{\text{Equiv}}$. Specifically, the matrix $\mathcal{A}_{\text{Equiv}}$ corresponding to the chosen subset of basis vectors is visualized inset. To clarify, the nodes $(S,v)$ that satisfy $v\in S$ “send messages” (i.e., broadcast their representation) to all the nodes $(S^{\prime},v^{\prime})$ such that $v=v^{\prime}$. In Section 5, we discuss in greater detail how this relates to GNN-SSWL$+$. A more formal discussion regarding our implementation of those symmetry based updates is given in Section F.4.
Maintaining sparsity. We note that while our characterization considers all nodes $(S,v)\in(\mathcal{P}([n])\times[n])$, in our specific choice of updates, we applied the updates only to the very sparse set of nodes in our product graph, i.e., $\mathcal{T}(G)\square G$. Nevertheless, this sparse update rule still preserves equivariance. This can be justified by applying a sparsity mask to exclude nodes not in the product graph before applying the linear equivariant layer over the full set of nodes. See discussion in [1].
Node Marking. While the bias term of the equivariant linear layer is the most natural choice to mark nodes in the product graph [26, 4, 12], our theoretical analysis in the next section reveals a more expressive variant, defined as follows,
$\mathcal{X}_{S,i}\leftarrow\sum_{j\in S}z_{\text{SPD}(i,j)}$ (10)
where $\text{SPD}(i,j)$ denotes the shortest path distance between nodes $i$ and $j$ in the original $G$ [^7] To facilitate this, we maintain a lookup table where each index corresponds to a shortest path distance, assigning a learnable embedding, $z_{\text{SPD}(i,j)}\in\mathbb{R}^{d}$, to each node $(S,i)$.. Note that replacing the sum in Equation 10 with any other invariant aggregator maintains equivariance.
Pooling. After a stacking of layers of the form Equation 2, we employ a pooling layer [^8] We note that for some of the theoretical analysis, we use: $\rho(\mathcal{X}^{\mathtt{T}})=\texttt{MLP}^{\mathtt{T}}\left(\sum_{S}\left(\texttt{MLP}^{\mathtt{T}}\big{(}\sum_{v=1}^{n}\mathcal{X}^{\mathtt{T}}(S,v)\big{)}\right)\right)$ to obtain a graph representation; switching back to the original notation (from indices $i$ to nodes $v$) that is, $\rho(\mathcal{X}^{\mathtt{T}})=\texttt{MLP}^{\mathtt{T}}\Big{(}\sum_{S}\Big{(}\sum_{v=1}^{n}\mathcal{X}^{\mathtt{T}}(S,v)\Big{)}\Big{)}$; $\mathtt{T}$ denotes the final layer.
5 Theoretical analysis
In this section, we study several theoretical properties of our framework. We start by analyzing different node marking strategies and then discuss the relation between our method, Subgraph GNNs, and coarsening-based approaches. Here, we present only the main results, further discussion is available in Appendices B, C and D, proofs in Appendix G.
Node marking strategies. The employed node-marking strategy described in Equation 10 constitutes a reasonable, yet specific design choice, but other compelling alternatives are possible. Here we suggest four natural marking variants and compare them in their expressiveness: Simple Node Marking $\pi_{S}$: element $(S,v)$ is ‘marked’ iff node $v$ is in super-node $S$; Node + Size Marking $\pi_{SS}$: element $(S,v)$ is ‘marked’ according to $\pi_{S}$ and, jointly, in a way to encode the size of $S$; Minimum Distance ($\pi_{MD}$): element $(S,v)$ is ‘marked’ with the minimum SPD between $v$ and the nodes in $S$; Learned Distance Function $\pi_{LD}$ (Equation 10): element $(S,v)$ is ‘marked’ by evaluating a (learned) permutation invariant function over the set of SPDs between $v$ and the nodes in $S$.
All the above variants represent natural extensions of the traditional marking techniques discussed in [4, 12, 27, 41], and constitute equally valid strategies for Higher-Order Subgraph GNNs operating on “unordered k-vertex subgraphs” [30, 27], which may be practically preferred to their ordered counterparts for efficiency reasons [^9] In practice, all these marking strategies can be implemented via Embedding Lookup layers.. We conveniently gather the first three strategies as $\Pi=\{\pi_{S},\pi_{SS},\pi_{MD}\}$ and summarize the relation between all variants as follows:
@@ 5 Theoretical analysis
(i) Strategies in $\Pi$ are all equally expressive, independently of the transformation function $\mathcal{T}$. Also, they are as expressive as uniquely encoding the elements’ orbit (as per Section 4.2.1, Lemma E.1). (ii) The strategy $\pi_{LD}$ is at least as expressive as strategies in $\Pi$. Additionally, there exists transformation functions s.t. it is strictly more expressive than all of them.
The above are formally stated in Propositions C.1, C.3 and C.2, and more thoroughly discussed in Appendix C. Importantly, they give theoretical justification to our specific choice of $\pi_{LD}$ (Equation 10). This is notably in contrast with the setting of node-based, full-bag Subgraph GNNs, where SPD-based techniques are generally as expressive as traditional node-marking [3, 39].
Recovering maximally expressive Subgraph GNNs An advantage of our approach is that the transformation function $\mathcal{T}$ can precisely control the behavior of our Subgraph GNN, allowing us to seamlessly interpolate between trivial message-passing equivalents on the original graph and maximally expressive node-based Subgraph GNNs, or to implement some higher-order variants thereof. The result below is formally stated in Proposition D.1 and further discussed in Section D.1:
Proposition 5.2 (Informal – Recovering the maximally expressive Subgraph GNNs).
For a trivial transformation function, namely $\mathcal{T}(G)=G$, CS-GNN can implement a maximally expressive Subgraph GNN as per [39].
More than a coarsening function. Our approach’s effectiveness comes not just from transforming the input graph ($\mathcal{T}(G)$), but from our carefully designed equivariant layers. These layers enable effective processing of the new product graph $\mathcal{T}(G)\square G$, capturing information from both $\mathcal{T}(G)$ and $G$ in a structured manner. We compare our method to a baseline that uses message-passing on the sum graph: the union of $\mathcal{T}(G)$ and $G$, with edges indicating node membership in super-nodes (defined formally in Definition D.2). We prove the following informal proposition, which we state formally in Propositions D.2 and D.3:
Proposition 5.3 (Informal – CS-GNN beyond coarsening).
@@ 6 Experiments
We experimented extensively over seven different datasets to answer the following questions: (Q1) Can CS-GNN outperform efficient Subgraph GNNs operating on small bags? (Q2) Does the additional symmetry-based updates boost performance? (Q3) Does CS-GNN in the full-bag setting validate its theory and match state-of-the-art full-bag Subgraph GNNs? (Q4) Does CS-GNN in the full-bag setting validate its theory and match state-of-the-art full-bag Subgraph GNNs?
@@
Peptides. We experimented on the Peptides-func and Peptides-struct datasets [9] – which full-bag Subgraph GNNs already struggle to process – evaluating CS-GNN’s ability to scale to larger graphs. The results are summarized in Table 2. By using an effective subset of subgraphs, CS-GNN outperforms all MPNN variants, even when incorporating structural encodings such as GATEDGCN+RWSE. Additionally, our method surpasses the random [^10] For the Peptides datasets, we benchmarked our model against the random variant of Subgraphormer + PE, which similarly incorporates information from the Laplacian eigenvectors. To ensure a fair comparison, we used a single vote. Additionally, since Subgraphormer + PE employs GAT [33] as the underlying MPNN, we also utilized GAT for this specific experiment to maintain consistency and fairness. baseline on both datasets.
@@
Ablation study – symmetry-based updates. We assessed the impact of the symmetry-based update on the performance of CS-GNN. Specifically, we asked, do the symmetry-based updates significantly contribute to the performance of CS-GNN? To evaluate this, we conducted several experiments using the Zinc-12k dataset across various bag sizes, $T\in\{2,3,4,5\}$, comparing CS-GNN with and without the symmetry-based update. The results are summarized in Table 3. It is clear that the symmetry-based updates play a key role in the performance of CS-GNN. For a bag size of $T=2$, the inclusion of the symmetry-based update improves the MAE by a significant $0.34$. For other bag sizes, the improvements range from $0.05$ to $0.016$, clearly demonstrating the benefits of including the symmetry-based updates.

### v4

@@ 4.1 Construction of the coarse product graph
$A^{\mathcal{T}}(S_{1},S_{2})=\begin{cases}1&\text{if }\exists v\in S_{1},\exists u\in S_{2}\text{ s.t. }A(v,u)=1,\\ 0&\text{otherwise},\end{cases}$ (3)
To clarify, in our running example (Figure 1), it holds that $A^{\mathcal{T}}(\{a,b,c,d\},\{e\})=1$, while $A^{\mathcal{T}}(\{a,b,c,d\},\{f\})=0$. For a more formal definition, refer to Definition A.3.
More specifically, our implementation of the graph coarsening function $\mathcal{T}$ employs spectral clustering [^6] Other graph coarsening or clustering algorithms can be readily used as well. [34] to partition the graph into $T$ clusters, which in our framework controls the size of the bag. This results in a coarsened graph with fewer nodes and edges than $G$. We highlight and stress that the space complexity of this sparse graph, $\mathcal{T}(G)$, is upper bounded by that of the original graph $G$ (we do not store $2^{n}$ nodes).
Defining the (coarse) product graph $\mathcal{T}(G)\square G$.
@@
The connectivity in this product graph induces the horizontal ($\mathcal{A}_{G}$) and vertical updates ($\mathcal{A}_{\mathcal{T}(G)}$) in Equation 2, visualized in Figure 1(right) via grey and red edges, respectively.
@@ 4.2 Symmetry-based updates
In the previous subsection, we used a combination of a coarsening function and the graph Cartesian product to derive the two induced connectivities $\mathcal{A}_{G},\mathcal{A}_{\mathcal{T}(G)}$ of our product graph. We use these connectivities to to perform message-passing on our product graph (see Equation 2).
Inspired by recent literature on Subgraph GNNs [12, 3, 39], which incorporates and analyzes additional non-local updates arising from various symmetries (e.g., updating a node’s representation via all nodes in its subgraphs), this section aims to identify potential new updates that can be utilized over our product graph. To that end, we study the symmetry structure of the node feature tensor in our product graph, $\mathcal{X}(S,v)$.The new updates described below will result in the third term in Equation 2, dubbed Symmetry-based updates ($\mathcal{A}_{\text{Equiv}}$). For better clarity in this derivation, we change the notation from nodes ($v$) to indices ($i$).
@@ 4.2.1 Symmetries of our product graph
where we define the action of $\sigma\in S_{n}$ on a set $S=\{i_{1},i_{2},\ldots,i_{k}\}$ of size $k$ as: $\sigma\cdot S\coloneq\{\sigma^{-1}(i_{1}),\sigma^{-1}(i_{2}),\ldots,\sigma^{-1}(i_{k})\}\coloneq\sigma^{-1}(S)$.
@@ 4.2.2 Derivation of linear equivariant layers for the node feature tensor
We now characterize the linear equivariant layers with respect to the symmetry defined above, focusing on Equation 6. We adopt a similar notation to [22], and assume for simplicity that the number of feature channels is $d=1$ (extension to multiple features is straightforward [22]). In addition, our analysis considers the case where $V^{\mathcal{T}}$ encompasses all potential super-nodes formed by subsets of $[n]$ (i.e we use the sparse coarsened adjacency [^8] This is because the action of $S_{n}$ is well defined over the index set $P(V[[n]])\times[n]$ but not over $V\times V^{\mathcal{T}}$).
Our main tool is the characterization of linear equivariant layers for permutation symmetries as parameter-sharing schemes [35, 31, 22]. In a nutshell, this characterization states that the parameter vectors of the biases, invariant layers, and equivariant layers can be expressed as a learned weighted sum of basis tensors, where the basis tensors are indicators of the orbits induced by the group action on the respective index spaces. We focus here on presenting the final results and summarize them in Proposition 4.1 at the end of this subsection. Detailed discussion and derivations are available in Appendix E.
Equivariant bias and invariant layers. The bias vectors of the linear layers in our space are in $\mathbb{R}^{2^{n}\times n}$. As shown in Figure 3(right), the set of orbits induced by the action of $S_{n}$ satisfies:
$(\mathcal{P}([n])\times[n])/S_{n}\coloneq\{\gamma^{k^{*}}:k=1,\ldots,n;*\in\{+,-\}\}.$ (7)
Here, $\gamma^{k^{+}}$ corresponds to all pairs $(S,i)\in\mathcal{P}([n])\times[n]$ with $|S|=k$ and $i\notin S$, and $\gamma^{k^{-}}$ to all pairs with $|S|=k$ and $i\in S$.
As stated in [35, 31, 22], the tensor set $\{\mathbf{B}^{\gamma}_{S,i}\}_{\gamma\in(\mathcal{P}([n])\times[n])/S_{n}}$ where:
@@
are a basis of the space of bias vectors of the invariant linear layers induced by the action of $S_{n}$.
@@
Weight matrices. Following similar reasoning, consider elements $(S_{1},i_{1},S_{2},i_{2})\in(\mathcal{P}([n])\times[n]\times\mathcal{P}([n])\times[n])$. In Appendix E we characterize the orbits of $S_{n}$ in this space as a partition in which each partition set is defined according to six conditions. Some of these conditions include the sizes of $S_{1}$, $S_{2}$ and $S_{1}\cap S_{2}$, which remain invariant under permutations. Given an orbit, $\Gamma\in(\mathcal{P}([n])\times[n]\times\mathcal{P}([n])\times[n])/S_{n}$, we define a basis tensor, $\mathbf{B}^{\Gamma}\in\mathbb{R}^{2^{n}\times n\times 2^{n}\times n}$ by setting:
@@
A visualization of the two basis vectors in Equations 8 and 9, is available in Figure 3. The following (informal) proposition summarizes the results in this section (the proof is given in Appendix G),
@@ 4.2.3 Incorporating symmetry-based updates in our framework
To begin, we note that any linear equivariant layer can be realized through an MPNN  [13] applied to a fully connected graph with appropriate edge features.
@@
This is formally stated in Lemma F.1, the main idea is to encode the parameters on the edges of this graph (see visualization inset). Thus, the natural construction of $\mathcal{A}_{\text{Equiv}}$ corresponds to a fully connected graph, with appropriate edge features derived from the parameter-sharing scheme we have developed.
@@
However, one of our main goals and guidelines in developing our flexible framework is to maintain efficiency, and to align with the (node-based) maximally expressive GNN, namely GNN-SSWL$+$ [39, 3], for the case of a trivial coarsening function, $\mathcal{T}(G)=G$ (which correspond to the full-bag setting). To achieve this, we opt for a sparser choice by using only a subset of the basis vectors (defined in Equation 9) to construct $\mathcal{A}_{\text{Equiv}}$. Specifically, the matrix $\mathcal{A}_{\text{Equiv}}$ corresponding to the chosen subset of basis vectors is visualized inset – the parameter-sharing scheme is represented by edges with matching colors. To clarify, the nodes $(S,v)$ that satisfy $v\in S$ “send messages” (i.e., broadcast their representation) to all the nodes $(S^{\prime},v^{\prime})$ such that $v=v^{\prime}$. A more formal discussion regarding our implementation of those symmetry based updates is given in Section F.4.
Maintaining sparsity. While the updates above are defined over the sparse representation of the coarse product graph, in practice we use its dense representation, treating it as a graph over the set of nodes $V\times V^{\mathcal{T}}$, which requires space complexity $\mathcal{O}(T\cdot|V|)$. The update rules above are adapted to this representation simply by masking all nodes $(S,v)$ in the sparse representation such that $S\notin V^{\mathcal{T}}$. We note the models using the resulting update rule remain invariant to the action of $S_{n}$. See discussion in [1].
4.3 Marking Strategies and Theoretical Analysis
One of the key components of subgraph architectures is their marking strategy. Two widely used approaches in node-based subgraph architectures are binary-based node marking [4] and distance-based marking [39], which were proven to be equally expressive in the full-bag setup [39]. Empirically, distance-based marking has been demonstrated to outperform other strategies across several standard benchmarks. In this section, our aim is to develop and theoretically justify an appropriate marking strategy, specifically tailored to the structure of our product graph. We present and discuss here our main results, and refer to Appendix C for a more formal discussion.
Building on existing marking strategies and considering the unique structure of our product graph, we propose two natural extensions to both the binary node marking [4] and distance-based marking strategies [39]. Extending binary node marking, we first suggest Simple Marking ($\pi_{S}$), where an element $(S,v)$ is assigned a binary feature that indicates whether node $v$ belongs to subgraph $S$ ($v\in S$). The second extension, Node + Size Marking ($\pi_{SS}$), builds on the simple marking by assigning an additional feature that encodes the size of the super-node $S$.
For distance-based strategies, we propose Minimum Distance ($\pi_{MD}$), where each element $(S,v)$ is assigned the smallest (minimal) shortest path distance (SPD) from node $v$ to any node $u\in S$. Finally, Learned Distance Function ($\pi_{LD}$) extends this further by assigning to each element $(S,v)$ the output of a permutation-invariant learned function, which takes the set of SPDs between node $v$ and the nodes in $S$ as input.
Surprisingly, unlike the node-based full-bag case, we find that these marking strategies are not all equally expressive. We conveniently gather the first three strategies as $\Pi=\{\pi_{S},\pi_{SS},\pi_{MD}\}$ and summarize the relation between all variants as follows:
@@ 4.3 Marking Strategies and Theoretical Analysis
(i) Strategies in $\Pi$ are all equally expressive, independently of the transformation function $\mathcal{T}$. (ii) The strategy $\pi_{LD}$ is at least as expressive as strategies in $\Pi$. Additionally, there exists transformation functions s.t. it is strictly more expressive than all of them.
The above is formally stated in Propositions C.1 and C.2, and more thoroughly discussed in Appendix C. In light of the above proposition, we instatiate the learned distance function $\pi_{LD}$ strategy when implementing our model, as follows,
$\mathcal{X}_{S,v}\leftarrow\sum_{u\in S}z_{d_{G}(v,u)}$ (10)
where $d_{G}(v,u)$ denotes the shortest path distance between nodes $v$ and $u$ in the original $G$ [^9] To facilitate this, we maintain a lookup table where each index corresponds to a shortest path distance, assigning a learnable embedding, $z_{d_{G}(v,u)}\in\mathbb{R}^{d}$, to each node $(S,v)$..
Coarsening Function and Expressivity. We investigate whether our CS-GNN framework offers more expressiveness compared to directly integrating information between the coarsened graph and the original graph.
The two propositions below illustrate that a simple, straight forward integration of the coarsen graph with the original graph (this integration is referred to as the sum graph – formally defined in Definition D.2), and further processing it via standard message-passing, results in a less expressive architecture. Furthermore, when certain coarsening functions are employed within the CS-GNN framework, our resulting architecture becomes strictly more expressive than conventional node-based subgraph GNNs. These results suggest that the interplay between the coarsening function and the subgraph layers we have developed enhances the model’s overall performance. We summarize this informally below and provide a more formal discussion in Appendix D.
Proposition 4.3 (Informal – CS-GNN goes beyond coarsening).
@@
Proposition 4.4 (Informal – CS-GNN vs node based subgraphs).
There exist transformations $\mathcal{T}$’s s.t. our CS-GNN model using $\mathcal{T}$ as its coarsening function is strictly more expressive than GNN-SSWL$+$.
@@ 5 Experiments
We experimented extensively over seven different datasets to answer the following questions: (Q1) Can CS-GNN outperform efficient Subgraph GNNs operating on small bags? (Q2) Does the additional symmetry-based updates boost performance? (Q3) Does CS-GNN offer a good solution in settings where full-bag Subgraph GNNs cannot be applied? (Q4) Does CS-GNN in the full-bag setting validate its theory and match state-of-the-art full-bag Subgraph GNNs?
@@
Peptides. We experimented on the Peptides-func and Peptides-struct datasets [9] – which full-bag Subgraph GNNs already struggle to process – evaluating CS-GNN’s ability to scale to larger graphs. The results are summarized in Table 2. CS-GNN outperforms all MPNN variants, even when incorporating structural encodings such as GATEDGCN+RWSE. Additionally, our method surpasses the random [^10] For the Peptides datasets, we benchmarked our model against the random variant of Subgraphormer + PE, which similarly incorporates information from the Laplacian eigenvectors. To ensure a fair comparison, we used a single vote and the same exact bag size of 35 subgraphs. Additionally, since Subgraphormer + PE employs GAT [33] as the underlying MPNN, we also utilized GAT for this specific experiment to maintain consistency and fairness. baseline on both datasets.
@@
Ablation study – symmetry-based updates. We assessed the impact of the symmetry-based update on the performance of CS-GNN. Specifically, we ask, do the symmetry-based updates significantly contribute to the performance of CS-GNN? To evaluate this, we conducted several experiments using the Zinc-12k dataset across various bag sizes, $T\in\{2,3,4,5\}$, comparing CS-GNN with and without the symmetry-based update. The results are summarized in Table 3. It is clear that the symmetry-based updates play a key role in the performance of CS-GNN. For a bag size of $T=2$, the inclusion of the symmetry-based update improves the MAE by a significant $0.034$. For other bag sizes, the improvements range from $0.005$ to $0.016$, clearly demonstrating the benefits of including the symmetry-based updates.